# Modular reduction of complex representations of finite reductive groups

### v1

@@ 6. On uniqueness of the elements $M_{w}$
In this section we explain in what sense our collection of elements $\{M_{w}\}_{w\in W}$ is unique in satisfying some of the properties conjectured in Conjecture 1.3. In particular, we show that if some collection $\{M_{w}^{\prime}\}_{w\in W}$ defined “in terms of $q$” (cf. Definition 6.5) satisfies Conjecture 1.1 and a precise version of property i) in Conjecture 1.3, then this collection is very close to our $\{M_{w}\}_{w\in W}$, in a sense which we make precise in Definition 6.1. The main result of this section is Corollary 6.7, which shows that with the aforementioned assumptions in hand, properties ii) and iv) of Conjecture 1.3 are impossible to satisfy.
@@
  • For all $w\in W$, $M_{w}^{\prime}$ is a linear combination of $V_{\lambda}$ for $\lambda$ being $\gamma$-close to $(q-1)\omega_{I(w)}$.
@@
We note that $A=\mathrm{ch}(V_{1,0}\otimes V_{0,1})$, and that $L_{1,0}=V_{1,0}$, $L_{0,1}=V_{0,1}$. By the Steinberg tensor product formula, we then note that $A=L_{p,1}=L_{1,p}$. By writing this as a linear combination of Weyl modules in two different ways (cf. [dynkin] for a computational method with which to do so), we see that
@@
Suppose that $G=SL(n)$ for $n\leq 5$. Then for any fixed $\gamma$, there exists $q^{\prime}>0$ such that if $q>q^{\prime}$ and $\{M_{w}^{\prime}\}_{w\in W}$ is a $\gamma$-close collection of candidate elements, then for all $w\in W$, $M_{w}^{\prime}-M_{w}$ is a linear combination of $V_{\lambda}$ for $\lambda$ which are $(\gamma+2\rho)$-close to 00.
@@
Let $\gamma$ be a fixed dominant weight. It is clear that we can choose $q^{\prime}$ large enough so that if $q>q^{\prime}$, then for any distinct subsets $I,I^{\prime}\subset\Pi$, there exist no weights $\lambda$ which are $(\gamma+2\rho)$-close to both $(q-1)\omega_{I}$ and $(q-1)\omega_{I^{\prime}}$.
@@
For any $w\in W$, $M_{w}^{\prime}-M_{w}$ is a linear combination of $V_{\lambda}$ for $\lambda$ being $\gamma$-close to $(p-1)\omega_{I(w)}$. It follows from classical modular representation theory that such a linear combination can then be written as a linear combination of $L_{\lambda}$ for $\lambda$ which is $(\gamma+\rho)$-close to $(q-1)\omega_{I(w)}$. By the Steinberg tensor product theorem, this can again be rewritten as a linear combination of $L_{\lambda}$ for $\lambda$ being $q$-restricted and $(\gamma+2\rho)$-close to either $(q-1)\omega_{I(w)}$ or to 00.
We now claim that in fact all $L_{\lambda}$ appearing in this linear combination with nonzero coefficient are of the second type, with $\lambda$ being $(\gamma+2\rho)$-close to 00. Suppose otherwise; then some $L_{\lambda}$ term for $\lambda$ which is $(\gamma+\rho)$-close to $(q-1)\omega_{I(w)}$ would have to also appear with nontrivial coefficient in the analogous expansion for $M_{y}^{\prime}-M_{y}$ for $y\neq w$, $y\in c$, so that these terms would cancel out in equation (40). But by our construction of this expansion, all of the $L_{\lambda}$ terms in the expression for $M_{y}^{\prime}-M_{y}$ must be $(\gamma+2\rho)$-close to either $(q-1)\omega_{I(y)}$ or to 00. By our choice of $q^{\prime}$, this means $\omega_{I(y)}=\omega_{I(w)}$ or $\omega_{I(y)}=0$. By Lemma 6.3, this implies $y=w$, which is a contradiction.
This means each $M_{w}^{\prime}-M_{w}$ is a linear combination of $L_{\lambda}$ (and therefore of $V_{\lambda}$) for $\lambda$ being $(\gamma+2\rho)$-close to 00. ∎
@@
Consider a family of collections of elements $\{M_{w}\}_{w\in W}$ defined simultaneously across all large values of $q$; we write $M_{w}^{q}$ when clarity is necessary to specify the dependence of the definition of $M_{w}$ on $q$. For a fixed $w\in W$, we say that $M_{w}$ is a linear combination of Weyl characters depending on $q$ if there exist polynomials $f_{i}(t)\in\mathbb{Z}[t][X^{*}(T)]$ such that
$M_{w}^{q}=\sum_{i}V_{f_{i}(q)}$
@@
We note that the failure of property iv), i.e. positivity, which is exhibited in the example given in Proposition 5.8, cannot be resolved by an adjustment by any linear combination of $V_{\lambda}$ for $\lambda$ close to 00, so long as $q$ is sufficiently large. So there exists no collection $\{M_{w}^{\prime}\}_{w\in W}$ satisfying the assumptions of the corollary such that $M_{s_{2}s_{3}s_{2}}^{\prime}$ satisfies posivitiy, since it must differ from $M_{s_{2}s_{3}s_{2}}$ by a linear combination of such $V_{\lambda}$.
@@
7. Comparison with exceptional collections
We now present a possible approach to understanding the constructions in the present paper on the level of coherent sheaves and derived categories on partial flag varieties $G/P$. We recall the definition of an exceptional collection and in Proposition 7.3 we explain how a $G$-equivariant exceptional collection on $G/P$ gives rise to an alternative construction of a lift of unipotent principal series representations from $G(\mathbb{F}_{q})$ to the algebraic group $G$. We then explain some evidence that for a particular exceptional collection (the one appearing in the paper [SVK]), such lifts agree with the ones guaranteed by the main construction in the present paper via our construction of the representations $\tilde{V}_{\chi}$. We conclude with a conjecture that this phenomenon might occur more generally.
@@ 7. Comparison with exceptional collections
Let $X$ be a projective variety over $\mathbb{F}_{q}$ with an action of the algebraic group $G$, then we can consider the representation $\mathsf{k}[X(\mathbb{F}_{q})]$. Suppose we have a decomposition
(41) $[\mathcal{O}_{\Delta}]=\sum_{i}[\mathcal{F}_{i}\boxtimes\mathcal{F}_{i}^{\prime}]$
@@
$[\mathcal{O}_{X({\mathbb{F}}_{q})}]=\sum[\mathcal{F}_{i}\otimes\mathrm{Fr}^{*}(\mathcal{F}_{i}^{\prime})]$
@@
We now recall the definition of an exceptional collection in a triangulated category $\mathcal{D}$ over a field. The only case we will consider is $\mathcal{D}=D^{b}(\mathrm{Coh}(G/P))$, where $P$ is a parabolic subgroup containing $B$ and $G/P$ is the corresponding partial flag variety over the field $\mathsf{k}$.
@@
An object $\mathcal{F}\in\mathcal{D}$ is exceptional if there is an isomorphism of graded $\mathsf{k}$-algebras $\mathrm{Hom}_{\mathcal{D}}^{\bullet}(\mathcal{F},\mathcal{F})\cong\mathsf{k}$. We say that a collection $(\mathcal{F}_{1},\dots,\mathcal{F}_{n})$ of exceptional objects is exceptional if for $1\leq i<j\leq n$ one has
$\mathrm{Hom}_{\mathcal{D}}^{\bullet}(\mathcal{F}_{j},\mathcal{F}_{i})=0.$
It is a full exceptional collection if the $\mathcal{F}_{i}$ generate $\mathcal{D}$ as a triangulated category.
A collection $(\mathcal{G}_{n},\dots,\mathcal{G}_{1})$ is said to be the dual exceptional collection to $(\mathcal{F}_{1},\dots,\mathcal{F}_{n})$ if
(43) $\mathrm{RHom}(\mathcal{F}_{i},\mathcal{G}_{j}[l])$ =$=\begin{cases}\mathsf{k}&i=j,l=0\\ 0&\textrm{otherwise},\end{cases}$
(44) $\mathrm{RHom}(\mathcal{G}_{i},\mathcal{F}_{j}[l])$ =$=\begin{cases}\mathsf{k}&i=j,l=0\\ 0&\textrm{otherwise}.\end{cases}$
In this case, the collection $(\mathcal{G}_{n},\dots,\mathcal{G}_{1})$ is also exceptional.
For any $\mathcal{F}\in D^{b}_{G}(G/P)$, we let $\mathcal{F}^{\vee}$ be its image under the Grothendieck duality functor $\mathcal{R}Hom(-,\mathcal{O}_{G/P})$, so that for any $\mathcal{G}\in D^{b}_{G}(G/P)$, $\mathrm{RHom}_{G/P}(\mathcal{F},\mathcal{G})=R\Gamma(G/P,\mathcal{F}^{\vee}\otimes\mathcal{G})$, where here the tensor product is left-derived.
@@
Suppose $C=\{\mathcal{F}_{w}\}$ is an exceptional collection generating the category $D^{b}(\mathrm{Coh}(G/P))$ and $\{\mathcal{F}^{w}\}$ is the dual collection, where the $\mathcal{F}_{w}$ and $\mathcal{F}^{w}$ are all $G$-equivariant. We define a virtual $\mathsf{k}$-representation of the algebraic group $G$ by
(45) $V_{C}$ $=\sum_{w}\mathrm{RHom}(\mathcal{F}_{w},\mathrm{Fr}^{*}(\mathcal{F}^{w}))$
Then the restriction of $V_{C}$ to $G(\mathbb{F}_{q})$ is the representation $\mathsf{k}[(G/P)(\mathbb{F}_{q})]$.
@@
First, we claim that
(46) $[\mathcal{O}_{\Delta}]=\sum_{w}[\mathcal{F}_{w}^{\vee}\boxtimes\mathcal{F}^{w}]$
in $K_{0}(\mathrm{Coh}^{G}((G/P)^{2}))$. Indeed, it is enough to check that the pairings of each side with any element of the form $[\mathcal{F}_{y_{1}}^{\vee}\boxtimes\mathcal{F}^{y_{2}}]$ under the pairing $([\mathcal{F}],[\mathcal{G}])\mapsto\chi(R\Gamma(\mathcal{F}\otimes\mathcal{G}))$ are equal. It is easy to check that both such pairings are equal to $\delta_{y_{1},y_{2}}$ by Grothendieck duality and the fact that $C$ is an exceptional collection.
By Lemma 7.1 for $X=G/P$, the result then follows from the fact that
(47) $\sum_{w}\mathrm{RHom}(\mathcal{F}_{w},\mathrm{Fr}^{*}(\mathcal{F}^{w}))=\sum_{w}R\Gamma(\mathcal{F}_{w}^{\vee}\otimes\mathrm{Fr}^{*}\mathcal{F}^{w})$
by Grothendieck duality. ∎
@@
Let $C_{X}^{P}=\{X_{w}\}_{w\in W}$ be the Samokhin-van der Kallen exceptional collection in $D^{b}(\mathrm{Coh}(G/P))$ constructed in [SVK]. We refer the reader to Section 9 of loc. cit. for details on this construction.

### v2

@@ 6. On uniqueness of the elements $M_{w}$
In this section we explain in what sense our collection of elements $\{M_{w}\}_{w\in W}$ is unique in satisfying some of the properties conjectured in Conjecture 1.3. In particular, we show that if some collection $\{M_{w}^{\prime}\}_{w\in W}$ defined “in terms of qq” (cf. Definition 6.5) satisfies Conjecture 1.1 and a precise version of property i) in Conjecture 1.3, then this collection is very close to our $\{M_{w}\}_{w\in W}$, in a sense which we make precise in Definition 6.1. The main result of this section is Corollary 6.7, which shows that with the aforementioned assumptions in hand, properties ii) and iv) of Conjecture 1.3 are impossible to satisfy.
@@
  • For all $w\in W$, $M_{w}^{\prime}$ is a linear combination of $V_{\lambda}$ for $\lambda$ being $\gamma$-close to $(q-1)\varpi_{I(w)}$.
@@
We note that $A=\mathrm{ch}(V_{1,0}\otimes V_{0,1})$, and that $L_{1,0}=V_{1,0}$, $L_{0,1}=V_{0,1}$. By the Steinberg tensor product formula, we then note that $A=L_{p,1}=L_{1,p}$. By writing this as a linear combination of Weyl modules in two different ways (cf. [undefe] for a computational method with which to do so), we see that
@@
Suppose that $G=SL(n)$ for $n\leq 5$. Then for any fixed $\gamma$, there exists $q^{\prime}>0$ such that if $q>q^{\prime}$ and $\{M_{w}^{\prime}\}_{w\in W}$ is a $\gamma$-close collection of candidate elements, then for all $w\in W$, $M_{w}^{\prime}-M_{w}$ is a linear combination of $V_{\lambda}$ for $\lambda$ which are $(\gamma+2\rho)$-close to 0.
@@
Let $\gamma$ be a fixed dominant weight. It is clear that we can choose $q^{\prime}$ large enough so that if $q>q^{\prime}$, then for any distinct subsets $I,I^{\prime}\subset\Pi$, there exist no weights $\lambda$ which are $(\gamma+2\rho)$-close to both $(q-1)\varpi_{I}$ and $(q-1)\varpi_{I^{\prime}}$.
@@
For any $w\in W$, $M_{w}^{\prime}-M_{w}$ is a linear combination of $V_{\lambda}$ for $\lambda$ being $\gamma$-close to $(q-1)\varpi_{I(w)}$. Such a linear combination can then be written as a linear combination of $L_{\lambda}$ for $\lambda$ which is $(\gamma+\rho)$-close to $(q-1)\varpi_{I(w)}$. By the Steinberg tensor product theorem, this can again be rewritten as a linear combination of $L_{\lambda}$ for $\lambda$ being $q$-restricted and $(\gamma+2\rho)$-close to either $(q-1)\varpi_{I(w)}$ or to 0.
We now claim that in fact all $L_{\lambda}$ appearing in this linear combination with nonzero coefficient are of the second type, with $\lambda$ being $(\gamma+2\rho)$-close to 0. Suppose otherwise; then some $L_{\lambda}$ term for $\lambda$ which is $(\gamma+\rho)$-close to $(q-1)\varpi_{I(w)}$ would have to also appear with nontrivial coefficient in the analogous expansion for $M_{y}^{\prime}-M_{y}$ for $y\neq w$, $y\in c$, so that these terms would cancel out in equation (40). But by our construction of this expansion, all of the $L_{\lambda}$ terms in the expression for $M_{y}^{\prime}-M_{y}$ must be $(\gamma+2\rho)$-close to either $(q-1)\varpi_{I(y)}$ or to 0. By our choice of $q^{\prime}$, this means $\varpi_{I(y)}=\varpi_{I(w)}$ or $\varpi_{I(y)}=0$. By Lemma 6.3, this implies $y=w$, which is a contradiction.
This means each $M_{w}^{\prime}-M_{w}$ is a linear combination of $L_{\lambda}$ (and therefore of $V_{\lambda}$) for $\lambda$ being $(\gamma+2\rho)$-close to 0. ∎
@@
Consider a family of collections of elements $\{M_{w}\}_{w\in W}$ defined simultaneously across all large values of $q$; we write $M_{w}^{q}$ when clarity is necessary to specify the dependence of the definition of $M_{w}$ on $q$. For a fixed $w\in W$, we say that $M_{w}$ is a linear combination of Weyl characters depending on $q$ if there exist polynomials $f_{i}(t)\in\mathbb{Z}[t][X^{*}(T)]$ and $c_{i}\in\mathbb{Z}$ such that
$M_{w}^{q}=\sum_{i}c_{i}V_{f_{i}(q)}$
@@
We note that the failure of property iv), i.e. positivity, which is exhibited in the example given in Proposition 5.8, cannot be resolved by an adjustment by any linear combination of $V_{\lambda}$ for $\lambda$ close to 0, so long as $q$ is sufficiently large. So there exists no collection $\{M_{w}^{\prime}\}_{w\in W}$ satisfying the assumptions of the corollary such that $M_{s_{2}s_{3}s_{2}}^{\prime}$ satisfies positivity, since it must differ from $M_{s_{2}s_{3}s_{2}}$ by a linear combination of such $V_{\lambda}$.
@@
7. Categorical interpretation by coherent sheaves on $G/P$
We now explain a perspective by which the constructions in the present paper can be understood in terms of coherent sheaves and derived categories on flag varieties. We define a collection of objects in $D^{b}(\mathrm{Coh}^{G}(T^{*}(G/P)))$ arising from the noncommutative Springer resolution, as in [undefa, undefb, undefc]. We explain how this collection gives rise to an alternative construction of a lift of the permutation representation $\mathsf{k}[(G/P)(\mathbb{F}_{q})]$ from $G(\mathbb{F}_{q})$ to the algebraic group $G$. We then show that this lift agrees with the main construction of the present paper, thereby giving a natural geometric interpretation for our construction.
@@ 7. Categorical interpretation by coherent sheaves on $G/P$
Let $X$ be a projective variety over $\mathbb{F}_{q}$ with an action of the algebraic group $G$, then we can consider the representation $\mathsf{k}[X(\mathbb{F}_{q})]$ of $G(\mathbb{F}_{q})$. Suppose we have a decomposition
(41) $\mathcal{O}_{\Delta}]=\sum_{i}[\mathcal{F}_{i}\boxtimes\mathcal{F}_{i}^{\prime}]$
@@
$\mathcal{O}_{X({\mathbb{F}}_{q})}]=\sum[\mathcal{F}_{i}\otimes\mathrm{Fr}^{*}(\mathcal{F}_{i}^{\prime})]$
@@
7.1. Comparison of bases and representations
We begin in the more standard case of $G/B$ for simplicity, and will then in §7.4 explain how to generalize the arguments presented in this subsection to the more general case of $G/P$. The objects we will introduce in this section arise in the study of the noncommutative Springer resolution and the constructible-coherent correspondence as in [undefa]; this correspondence and these objects are explained in [undefb] and [undefc]. The results in loc. cit. establish a derived equivalence
(43) $D^{b}(\mathrm{Coh}^{G}(T^{*}(G/B)))\cong D^{b}(\mathsf{A}\mathrm{-mod}^{G})$
for an algebra $\mathsf{A}$ with a $G$-action. The categories in (43) can then also be identified with the derived category of Iwahori-Whittaker sheaves on the affine flag variety, which is the main result of [undefa]. The image of the tautological $t$-structure on the right-hand side of (43) in $D^{b}(\mathrm{Coh}^{G}(T^{*}(G/B)))$ is called the noncommutative Springer $t$-structure.
We note that the bases considered in this section are closely related to the Kazhdan-Lusztig basis for the affine Hecke algebra and its spherical quotient, along with the canonical basis for Lusztig’s periodic Hecke module as introduced in [undeft, undefo]. Thus we will make use of the extended affine Weyl group $W^{\mathrm{aff}}$ and its Kazhdan-Lusztig basis $\{C_{w}^{\prime}\}_{w\in W^{\mathrm{aff}}}$. We will also use the natural inclusion $t:X^{*}(T)\to W^{\mathrm{aff}}$.
For the remainder of this section, we assume that $G$ is simply connected, which guarantees the $G$-equivariance in the following definition.
@@ 7.1. Comparison of bases and representations
Let $\{E_{i}\}$ (resp. $\{E^{i}\}$) denote the collection of irreducible (resp. projective) modules over the algebra $\mathsf{A}$. These modules are constructed explicitly in [undefc] and each is shown to carry a $G$-equivariant structure. Thus we let $\{\mathcal{E}_{i}\}$ and $\{\mathcal{E}^{i}\}$ denote their images in $D^{b}(\mathrm{Coh}^{G}(T^{*}(G/B)))$ under the equivalence in (43).
Definition 7.3.
For any $w\in W$, let
(44) $g_{w}$ $=\frac{1}{|W_{I(w)}|}C_{w\cdot t(\varpi_{\overline{I(w)}})}^{\prime}\cdot 0\in N,$
where $C_{y}^{\prime}$ is the affine Kazhdan-Lusztig basis element corresponding to $y\in W^{\mathrm{aff}}$.
The next proposition will use the natural identification $K_{0}(\mathrm{Coh}^{G}(T^{*}(G/B)))\cong K_{0}(\mathrm{Coh}^{G}(G/B))\cong\mathbb{Z}[T]$. We continue to use the natural action of $\mathbb{Q}[W]$ on $\mathbb{Q}[T]$.
@@
The collections $\{\mathcal{E}_{i}\}$ and $\{\mathcal{E}^{i}\}$ have a natural indexing by the Weyl group, and thus can be written $\{\mathcal{E}_{w}\}_{w\in W}$, $\{\mathcal{E}^{w}\}_{w\in W}$.
Further, under the isomorphism
$K_{0}(\mathrm{Coh}^{G}(T^{*}(G/B)))\cong N,$
the image of $\mathcal{E}_{w}]$ is $g_{w}$, while the image of $\mathcal{E}^{w}]$ is the basis vector dual to $g_{w}$ with respect to the Euler pairing.
@@
First, note that there is a natural map
(45) $D^{b}(\mathrm{Coh}^{G}(T^{*}(G/B)))\to D^{b}(\mathrm{Coh}^{T}(T^{*}(G/B)))$
obtained by forgetting $G$-equivariance. For $\mathcal{F}\in D^{b}(\mathrm{Coh}^{G}(T^{*}(G/B)))$, we denote by $\mathcal{F}^{T}$ its image in $D^{b}(\mathrm{Coh}^{T}(T^{*}(G/B)))$ under this map.
Recall that $K_{0}(\mathrm{Coh}^{T}(G/B))$ can naturally be identified with Lusztig’s periodic Hecke module as defined in [undeft, undefo]. As explained in loc. cit. and [undefm], the periodic Hecke module has a canonical basis $\{\tilde{C}_{w}^{\prime}\}_{w\in W^{\mathrm{aff}}}$.
First we claim that elements of the form $\mathcal{E}_{i}^{T}(\lambda)]\in K_{0}(\mathrm{Coh}^{T}(T^{*}(G/B)))$ across all $i$ and across all $\lambda\in X^{*}(T)$ are exactly canonical basis elements of the form $\tilde{C}_{w}^{\prime}$ for $w\in W^{\mathrm{aff}}$. By their construction the objects $\mathcal{E}_{i}^{T}$ are each set-theoretically supported on the zero section, and since $\mathrm{pr}_{*}:K_{0}(\mathrm{Coh}^{T}(T^{*}(G/B)))\to K_{0}(\mathrm{Coh}^{T}(G/B))$ is an isomorphism when restricted to such objects, it suffices to show this for the objects $\mathrm{pr}_{*}(\mathcal{E}_{i}^{T}(\lambda))]\in K_{0}(\mathrm{Coh}^{T}(G/B))$. This is exactly what is shown in [undefc, §5.3.6] (for $e=0$). The main result of [undefm] ensures that the canonical bases appearing in this theorem (which are defined from a $K$-theoretic perspective in [undefl]) agree with the basis $\tilde{C}_{w}^{\prime}$ defined from a combinatorial perspective in [undeft, undefo]. Similarly, each $(\mathcal{E}^{i})^{T}(\lambda)]$ is a dual basis element under the natural pairing on the periodic Hecke module defined combinatorially in [undefo] and $K$-theoretically in [undefl], c.f. [undefc, §5.3.6].
Since this shows that for every $i$, the elements $\mathcal{E}_{i}^{T}(\lambda)]$ with $\lambda\in X^{*}(T)$ form a full $X^{*}(T)$-orbit of basis elements, we obtain a bijection from the set of indices $i$ to the set $W^{\mathrm{aff}}/X^{*}(T)\cong W$. Thus for any $w\in W$, we have characterized $\mathcal{E}_{w}^{T}$ as the unique object in the image of the forgetful functor (45) such that $\{[\mathcal{E}_{w}^{T}(\lambda)]\}_{\lambda\in X^{*}(T)}=\{\tilde{C}_{w\cdot t(\mu)}^{\prime}\}_{\mu\in X^{*}(T)}$.
Further, it follows from the identifications of [undefc, undefm] that $\mathcal{E}_{w}^{T}]=\frac{1}{|\mathrm{stab}_{W}(\mu)|}\tilde{C}_{w\cdot t(\mu)}^{\prime}$ for the unique $\mu\in X^{*}(T)$ satisfying $\langle\mu,\alpha_{i}^{\vee}\rangle=0$ if $w(\alpha_{i})$ is a positive root and $\langle\mu,\alpha_{i}^{\vee}\rangle=1$ if $w(\alpha_{i})$ is a negative root. We note that $\mu=\varpi_{\overline{I(w)}}$ satisfies these conditions, and thus for any $w\in W$ we have
$\mathcal{E}_{w}^{T}]=\frac{1}{|W_{I(w)}|}\tilde{C}_{w\cdot t(\varpi_{\overline{I(w)}})}^{\prime}.$
Under the identification between $\mathbb{C}[T]$ and $K_{0}(\mathrm{Coh}^{G}(T^{*}(G/B)))$, and between the periodic module and $K_{0}(\mathrm{Coh}^{T}(T^{*}(G/B)))$, the forgetful map in (45) sends $C_{w\cdot t(\varpi_{\overline{I(w)}})}^{\prime}\cdot 0\in\mathbb{C}[T]$ to $\tilde{C}_{w\cdot t(\varpi_{\overline{I(w)}})}^{\prime}$ (see equation (54) and the subsequent remark in [undef, §6.1.8]).
The statement about $\mathcal{E}^{w}]$ follows from the same discussion together with the compatibility of the natural Euler pairings under the above identifications. ∎
For any smooth $G$-variety $X$ and $\mathcal{F}\in D^{b}(\mathrm{Coh}^{G}(X))$, we let $\mathcal{F}^{\vee}$ be its image under the Grothendieck duality functor $\mathcal{R}Hom_{X}(-,\mathcal{O}_{X})$. In particular, for any $\mathcal{F},\mathcal{G}\in D^{b}(\mathrm{Coh}^{G}(G/B))$, we have
$\mathrm{RHom}_{G/B}(\mathcal{F},\mathcal{G})=R\Gamma(G/B,\mathcal{F}^{\vee}\otimes\mathcal{G}),$
where the tensor product is left-derived.
In the following, we let $i:G/B\to T^{*}(G/B)$ be the inclusion of the zero section, and $\mathrm{pr}:T^{*}(G/B)\to G/B$ be the projection.
@@
We define a virtual $\mathsf{k}$-representation of the algebraic group $G$ by
(46) $\tilde{V}_{\mathcal{E}}$ $=\sum_{w}\mathrm{RHom}(i^{*}\mathcal{E}^{w},\mathrm{Fr}^{*}(\mathrm{pr}_{*}\mathcal{E}_{w})).$
We now state the main theorem of this section.
Theorem 7.6.
Let $V=\mathsf{k}[(G/B)(\mathbb{F}_{q})]$. Then the virtual $\mathsf{k}$-representations $\tilde{V}$ (constructed in Section 2.3 using the basis $\{f_{w}\}_{w\in W}$) and $\tilde{V}_{\mathcal{E}}$ (constructed in Definition 7.5, see also Proposition 7.9) are equal.
Moreover, if $w\in\mathcal{J}$, $d$ is the Duflo involution in the same left cell as $w$, and
(47) $\mathcal{M}_{w}$ $=\mathrm{RHom}(i^{*}\mathcal{E}^{w_{0}w},\mathrm{Fr}^{*}(\mathrm{pr}_{*}\mathcal{E}_{w_{0}d})),$
then $M_{w}=[\mathcal{M}_{w}]$ in the Grothendieck group of virtual $\mathsf{k}$-representations of $G$.
To prove Theorem 7.6, we now introduce an intermediate object $\tilde{V}_{\mathrm{KL}}$, which we will then compare to both $\tilde{V}$ and $\tilde{V}_{\mathcal{E}}$. Recall that in Section 2.3, we explained that while the lift $\tilde{V}$ of $V$ to a representation of the algebraic group $G$ over $\mathsf{k}$ we constructed used the specific choice of basis $f_{w}$ for the $A$-module NN, such a lift is still well-defined for any $A$-basis for $N$ satisfying the axioms in Lemma 3.1. As such, we can make the following definition.
Definition 7.7.
Let $\tilde{V}^{\mathrm{KL}}$ be the lift of $V$ arising by the construction of Section 2.3 from the basis $\{g_{w}\}_{w\in W}$. This collection is a basis for $N$ as an $A$-module.
To prove Theorem 7.6, we first show that $\tilde{V}_{\mathcal{E}}\cong\tilde{V}_{\mathrm{KL}}$, then $\tilde{V}_{\mathrm{KL}}\cong\tilde{V}$.
Proposition 7.8.
The virtual representations $\tilde{V}^{\mathrm{KL}}$ and $\tilde{V}_{\mathcal{E}}$ are equal.
Proof.
We have
(48) $\tilde{V}_{\mathcal{E}}=\sum_{w\in W}R\Gamma_{G/B}(i^{*}(\mathcal{E}^{w})^{\vee}\otimes\mathrm{Fr}^{*}(\mathrm{pr}_{*}\mathcal{E}_{w})).$
By Proposition 7.4, the classes of $\{\mathcal{E}_{w}\}_{w\in W}$ identify with the basis $\{g_{w}\}_{w\in W}$, while the classes of $\{\mathcal{E}^{w}\}_{w\in W}$ identify with the dual basis. Thus (48) is exactly the expression for the lift $\tilde{V}^{\mathrm{KL}}$ arising from the basis $\{g_{w}\}_{w\in W}$. ∎
Proposition 7.9.
The restriction of $\tilde{V}_{\mathcal{E}}$ to $G(\mathbb{F}_{q})$ is the representation $\mathsf{k}[(G/B)(\mathbb{F}_{q})]$.
Proof.
From (48), we have
(49) $\tilde{V}_{\mathcal{E}}=\sum_{w\in W}R\Gamma_{G/B}(i^{*}(\mathcal{E}^{w})^{\vee}\otimes\mathrm{Fr}^{*}(\mathrm{pr}_{*}\mathcal{E}_{w})).$
Thus it remains to show that
(50) $[\mathcal{O}_{\Delta}]=\sum_{w}[(\mathcal{E}^{w})^{\vee}\boxtimes\mathcal{E}_{w}]$
in $K_{0}(\mathrm{Coh}^{G}(T^{*}(G/B)^{2}))$. By the equivalence to $D^{b}(\mathsf{A}\mathrm{-mod}^{G})$ established in [undefa, undefc], this reduces to checking
(51) $[\mathsf{A}_{\mathrm{diag}}]=\sum_{w}[(E^{w})^{\vee}\boxtimes E_{w}],$
where $\mathsf{A}_{\mathrm{diag}}$ is the diagonal $\mathsf{A}$-bimodule. This follows from the fact that $\{E^{w}\}_{w\in W}$ are the indecomposable projective $\mathsf{A}$-modules and $\{E_{w}\}_{w\in W}$ are the corresponding irreducible modules. ∎